The data in the table to the right give the total cost, C, and marginal cost, MC, for a firm at each possible level of outputFurther, profit will from $to $
Answer #1

P=62

it means MR=$62

as the price of the market is given means, the firm is in a perfectly competitive market where

MR=P=AR

P=$52 MR=$52

decrease from 11 units to 10units The firm chooses output at MC=P or nearest

above MC

at P=62

Q=11

at P=52

Q=10 units

profit will decrease from $150 to $26

profit=TR-TC

TR=P*Q

at P=62

TR=62*10=620

Profit=620-475=$145

at P=52

TR=52*10=520

Profit=520-410=$110

1)decrease

2)decrease